Circular No. 7177

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA VELORUM 1999 G. J. Garradd, Loomberah, N.S.W., reports the following accurate

CCD position obtained with a 0.45-m f/5.4 reflector on May 23.358 UT: α = 10^{h}44^{m}48^{s}.39, δ = −52^{o}25′30′′.7 (equinox 2000.0; 6 ACT stars, rms residuals 0′′.29). He notes a distinct pale yellow color of the nova visually through the same telescope. P. Schmeer, Bischmisheim, Germany, notes that there is a blue star (mag 16.4) listed in the USNO-A2.0 catalogue at position end figures 48^{s}.416, 31′′.23. Visual magnitude estimates: May 23.333, 2.6 (D. A. J. Seargent, The Entrance, N.S.W.); 23.508, 2.6 (A. Pearce, Nedlands, W. Australia).

η CARINAE

M. F. Corcoran and S. A. Drake, Universities Space Research Association, Goddard Space Flight Center (GSFC); J. H. Swank and R. Petre, Laboratory for High Energy Astrophysics, GSFC; K. Ishibashi, K. Davidson and R. Humphreys, University of Minnesota; A. Damineli, Instituto Astronômico e Geofísico da University de São Paulo; K. Koyama and Y. Tsuboi, Kyoto University; and R. Viotti, Instituto Astrofisica Spaziale, CNR, Rome, report: “We compared two x-ray spectra of the peculiar variable star η Car obtained by the ASCA observatory on 1993 Aug. 24 and 1999 Feb. 2. These spectra were both obtained near phase 0.20, where phase is calculated according to the epoch and period of Damineli et al. (1997, New Astron. 2, 107). We find that the absorbed 2–10-keV x-ray flux in 1999 is (6.4 ± 0.05) × 10^{−11} erg cm^{−2} s^{−1}, while that in 1993 was (5.8 ± 0.09) × 10^{−11} erg cm^{−2} s^{−1}. Our RXTE monitoring of the 2–10-keV x-ray flux for ± 20 days around the ASCA observation (1999 Jan. 20–Mar. 1) shows a steady decline in this energy band (from 23 to 18 PCU counts/s), with no sign of any x-ray flaring activity. Thus the observed change in the ASCA spectra suggests that the 2–10-keV x-ray flux has undergone a long-term brightening of ∼ 10 percent that is independent of the Damineli cycle. The x-ray variation may be related to a long-term increase in the optical brightness of the star (cf. van Genderen et al. 1999, A.Ap. 343, 847) and/or the recent brightening at optical and infrared wavelengths (IAUC 7146). The x-ray brightening we report is not the result of a change in circumstellar absorption, since circumstellar absorption is negligible at energies above 2 keV. Our preliminary analysis suggests that the flux increase is primarily due to an increase in the emission measure of the x-ray emitting region during the interval 1993–1999.”

1999 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green